Rapheal paid $1,148 for a camera during a 30% off sale. What was the cameras regular price.

Answer by jim_thompson5910(35256)

x = original price

30% of x = 0.30x (this is the amount you save, aka the discount)

-------------------------------------------------------

x - 0.30x = 0.70x

So the final price is 0.70x, or 70% of the original price

0.70x = 1148

0.70x/0.70 = 1148/0.70

x = 1640


The camera's original price is $1,640 (this is before the discount is applied)